OpenStudy (anonymous):

Which chemical equation is balanced? (1 point) H20 right arrow H2 + 2O2 2H20 right arrow 2H2 + O2 2H2O right arrow H2 + 2O2 H20 right arrow 2H2 + 2O2

OpenStudy (anonymous):

2H2O ---> 2H2 + O2

OpenStudy (anonymous):

2H2= 4 hydrogen O2= 2 oxygen SO 2H2O= H2*2=4 H O*2=2O this matches the above amount of hydrogen and oxygen

OpenStudy (anonymous):

Connections?